2 4. Solve the initial value problem: X 1 1 4 0 2 0 1 1 1 X X(0) = 6 0
Answer #1

4) The given initial value problem is xa 1 1 O 0 2 X with X(0) - Nuevo No - Luxe have to soive the system: The vou can be so

In the given problem, A = 1 1 4 a 0 2 — hoxe to corresponding First determine the rigen vawes and eigenwectors of A To find t

AV, = 1,2 1 4 0 - 2 2 :)) B - o, +1 +471 221 2B1 2B1 di+ßitri 271 di+ß,+47, 20. - ditBitri=bri -0, +Bi +41,=0 ar dit ßi-8=0 B

Hence U = - 3 2 Let = to the ** (3) Com responding be the eigenvector BL Y eigenvalue 12:-) suih that, A = A₂ U2 - * C ;))--

d) 2₂ + ß3 + 483 383 283 3 ß3 az + ß3 +²3 37 ar dz + B3 +4 %3 = 303 3 233 = 363 03+B3+ Y3-373 ~ 38, -233 = 0 - 203 +B3+4 Y3 -

GY, 54-26 +26, -2 -30=6 and 24+2+ (z=0 ~, 5(-2)-244 +243 = 2 C = -2 OY, 24-2)+(2+63 =0 - 262 +263 = 2+10 on C₂ + (3 =4 -C2 +

The solution of the given initial value problem (i) is X(+) 11 -1002t + 24-4710834 Geet -402t-e-t +503 +